OpenStudy (anonymous):

Describe pattern show in each graph. (-4,-4), (-2,-2), (0,0) (2,2_

OpenStudy (anonymous):

graph missing

OpenStudy (anonymous):

You don't need the graph to describe the pattern. I gave you cordinates. ;3

OpenStudy (anonymous):

approximate pattern for triangle

OpenStudy (anonymous):

Not that triangle the pattern it does to get to next number

OpenStudy (anonymous):

addition to each (x,y ) coordinate by 2

OpenStudy (anonymous):

example -4+2 = -2 -2+2=0 0+2=2 2+2=4

OpenStudy (anonymous):

Looks pretty linear to me....

OpenStudy (anonymous):

same is repeated with y

OpenStudy (anonymous):

linear, as in a line....